Antonio and Jeremy work for a plumbing company that was recently awarded a contract to install the plumbing fixtures in a new office complex. There are 12 bathrooms in the building. It takes Jeremy 4 hr longer than Antonio to complete one bathroom. If they work together it takes 8 hr to complete a bathroom. Find the rate at which each individual can complete a bathroom working alone.
Found 2 solutions by jorel555, ikleyn:
Answer by jorel555(1290)

Let 1/n be the amount of work Antonio can do in one hour. Then:
1/n+1/(n+4)=1/8
8/n+8/(n+4)=1
8(n+4)+8n=nē+4n
nē-12n-32=0
Using the quadratic formula, we get n=14.246 hrs, and n+4=18.246 hrs. ☺☺☺☺

Answer by ikleyn(52788)

.
Antonio and Jeremy work for a plumbing company that was recently awarded a contract to install the plumbing fixtures
in a new office complex. There are 12 bathrooms in the building. It takes Jeremy 4 hr longer than Antonio to complete one bathroom.
If they work together it takes 8 hr to complete a bathroom.
Find the rate at which each individual can complete a bathroom working alone.
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~

There are many other solved problems on joint work. See them in the lessons
    - Using Fractions to solve word problems on joint work
    - Solving more complicated word problems on joint work
    - Using quadratic equations to solve word problems on joint work
    - Selected joint-work word problems from the archive